# biblio.bib

### v1

@@ 1. Introduction and main result
In this paper we provide the first rigorous justification of the “gnocchi phase”, which is one of the nuclear pasta phases that are believed to form in the outer crust of neutron stars. We work with the simplest physical model used to describe this situation, called the liquid drop. It is due to Gamow [Gamow-29, Gamow-30] and was later refined by Bohr and Wheeler [BohWhe-39]. The same model can also describe microphase separation of diblock copolymers [OhtKaw-86].
@@
Here ${\rm Per}\,(\Omega)$ denotes the perimeter of the set $\Omega$ in the sense of geometric measure theory, see [Maggi-12] for a textbook introduction. When $\Omega$ has a smooth (in fact, Lipschitz is enough) boundary, then ${\rm Per}\,(\Omega)$ coincides with its ordinary surface area. The perimeter takes into account the short range interactions between the nuclei in a droplet. The second term is the electrostatic energy due to the fact that electrons and protons are, respectively, negatively and positively charged. The constant $\rho$ is the relative fraction of delocalized electrons. We have chosen units so that the local nuclear density and the constant in front of the perimeter are both equal to 1. In physical units they depend on the proton-neutron ratio. Note that protons are not able to bind without neutrons.
@@
Since $\Omega\subset\Lambda$, this implies $0\leqslant\rho\leqslant 1$ and then by symmetry we can assume $20\leqslant\rho\leqslant 1/$ 2.
At equilibrium the nuclear phase $\Omega\subset\Lambda$ minimizes the energy given in (3). Working under the constraint that the system is exactly neutral, the minimum equals
@@
Our goal in this paper is to study the shape of these minimizers in the low density regime. More precisely, we will look at the resulting energy. We will first take $\Lambda\nearrow\mathbb{R}^{3}$ before we actually study the limit $\rho\to 0$. This order of the limits is important for the result to be physically meaningful.
@@ 1.1. Thermodynamic limit
Let $0\leqslant\rho\leqslant 1$. Let $\Lambda_{n}$ be a sequence of $(r,L)$–Lipschitz open sets, so that
@@
For us an open set $\Lambda\subset\mathbb{R}^{3}$ is $(r,L)$–Lipschitz if for any $x\in\partial\Lambda$, the set $\Lambda\cap B(x,r)$ is, in an appropriate coordinate system, equal to the epigraph of an $L$–Lipschitz function. Under this assumption, one can see that the Fisher condition (5) holds for $h\leqslant Cr$. Hence (5) is only an assumption for large $h$. Note that all the conditions are satisfied if $\Lambda_{n}$ is the rescaling of a fixed $(r,L)$–Lipschitz bounded open set containing the origin (for instance a cube or a ball). Our assumptions on the domain $\Lambda_{n}$ are not sharp and have been chosen because they are relatively simple.
@@
The existence of the limit (7) for the periodic model was shown first in [AlbChoOtt-09] who also provided a convergence rate and information about the minimizers. That this also implies that the limit in (6) exists for cubes and coincides with (7) was noted in [EmmFraKon-20]. In Appendix A we provide a completely different proof of Theorem 1 and of the limit (7) using classical methods for Coulomb systems introduced by Lieb, Lebowitz and Narnhofer in the 1970s [LieLeb-72, LieNar-75]. This proof works for any sequence $\Lambda_{n}$ satisfying the mentioned conditions, hence it is much more flexible than the existing approaches.
@@
The next term in the expansion (8) describes how the gnocchis are arranged in the tomato sauce, corresponding to the low-density uniform background of electrons. The Jellium model (introduced below) tells us that the gnocchis will be placed at large distances, of order $\rho^{-1/3}$. A famous conjecture of Wigner [Wigner-34] states that they will in fact be regularly arranged on a Body-Centered Cubic lattice, in which case the Jellium energy is $e_{\rm Jel}\simeq-1.4442$. The interpretation of the negativity of $e_{\rm Jel}$ is that there is an effective attraction between the droplets, mediated by the background of electrons.
Figure 1. Illustration of Theorem 2 in the two-dimensional case. For small background density $\rho$, a minimizer $\Omega$ will consist of many droplets of size one and relative distances of order $\rho^{-1/3}$.
@@
Let us mention some previous works related to Theorem 2. The first term $\mu_{*}\rho$ was derived in [KnuMurNov-16] in a very dilute regime where the limits $L\to\infty$ and $\rho\to 0$ are coupled in the manner $\rho\sim 1/L^{2}$. The leading order in the regime considered here (first $L\to\infty$ and then $\rho\to 0$) was derived in [EmmFraKon-20]. The ultra-dilute regime with coupled limits $\rho\sim 1/L^{3}$ was considered earlier in [ChoPel-10] for the periodic model and a two-term expansion in the same spirit as (8) was proved. This is a regime where screening effects are much less pronounced than in the situation considered in this paper.
@@
The main argument we use for proving (8) is to reduce to smaller sets of size $A\rho^{-1/3}$ with $A$ large but fixed. In any such set we expect finitely many droplets, similar to the regime studied in [ChoPel-10]. For the upper bound we involve the periodic Jellium problem defined as in (7) but for point-like particles. For the lower bound we use the Graf–Schenker inequality as in [GraSch-95, HaiLewSol_thermo-09] to remove the interactions between the different sets. We then argue by compactness to expand the energy locally in the corresponding sets of size $A\rho^{-1/3}$.
Several phase transitions are believed to occur when the density is increased from 00 to $1/2$, that is, when going deeper into the crust of the neutron star. Those are often called “pasta phases” [BohWhe-39, RavPetWil-83, HasSekYam-84, Oyamatsu-93]. At small but positive $\rho$ the gnocchis are not believed to be perfect balls as they become in the limit $\rho\to 0$. They will be slightly elongated with an asymmetry increasing with $\rho$. When the density is increased beyond a certain value, instead of gnocchis one expects infinitely long cylinders called “spaghetti”, which are arranged along a two-dimensional lattice. (Just like the gnocchis may not be perfect balls, the base of the spaghetti cylinders may not be perfect disks.) When the density is further increased beyond another critical value (strictly less than $1/2$), instead of spaghettis one expects infinite slabs called “lasagna”, which are arranged along a one-dimensional lattice. In particular, at $\rho=1/2$ nucleons are expected to occupy slabs that are separated by completely unoccupied slabs of the same thickness. When $\rho$ is further increased, the story goes in the reverse way for the complement $\mathbb{R}^{d}\setminus\Omega$ until we reach an anti-gnocchi phase with $\Omega$ occupying most of $\mathbb{R}^{3}$, with holes looking likes balls at distances $(1-\rho)^{-1/3}$ and an energy behaving like for $\rho\to 0$, by symmetry. The energies in the spaghetti and lasagna regimes are given by the one- and two-dimensional analogues of the model studied here. The one-dimensional analogue of $e(\rho)$ can be computed explicitly [RenWei-00, FraLie-19]. The two-dimensional problem is still open and for partial results we refer to [CheOsh-07, GolMurSer-13, GolMurSer-14, Emmert-25_PhD]
@@ 1.3. Isolated droplets
$\boxed{\mu_{*}:=\min_{\begin{subarray}{c}\Omega\subset\mathbb{R}^{3}\\ |\Omega|>0\end{subarray}}I[\Omega],\qquad I[\Omega]:=|\Omega|^{-1}\left({\rm Per}\,\Omega+\frac{1}{2}\iint_{\Omega^{2}}\frac{\mathrm{d}x\,\mathrm{d}y}{|x-y|}\right).}$ (9)
@@
It is known [KnuMur-13, FraLie-15, KnuMurNov-16] that the minimum in (9) is attained for some bounded set $\Omega_{*}$, and that the problem admits no minimizer of a too large volume or too large diameter (but the infimum can be the same because we can put several copies of $\Omega_{*}$ infinitely far apart). This allows us to introduce the largest possible volume of minimizers:
@@
It is rigorously known that $m_{*}\leqslant 8$ by [FraKilNam-16] and $m_{**}\geqslant 5/2$ by [FraLie-15, Prop. 4.1].
For our proof of Theorem 2, we will need to know the rather precise behavior of minimizing sequences for the minimization problem $\mu_{*}$ in (9). The following result is an improvement of [FraLie-15, KnuMurNov-16].
@@ 1.4. Jellium model
Jellium is often associated with the name of Wigner [Wigner-34] who however introduced the quantum version in 1934. The classical model is also called the one-component plasma and it seems to have been proposed around 1900 by J. J. Thomson [Thomson-04] – based on previous ideas of W. Thomson (Lord Kelvin) – in order to describe the electrons in an atom, before the discovery of the nuclei. In this context, it is sometimes called the plum pudding model.
@@
The existence of the limit (16) was proved first by Lieb and Narnhofer in 1975 [LieNar-75], based on a method introduced earlier by Lieb and Lebowitz in [LieLeb-72]. This model has recently reappeared in many works mentioned in the review paper [Lewin-22]. In particular, in a long series of works culminating in [Serfaty-24_lecture_notes] Serfaty and co-workers have obtained detailed results about the Jellium model with variable background density. As a byproduct, also results about the original Jellium model were obtained.
@@
$-1.4508\approx-\frac{3}{5}\left(\frac{9\pi}{2}\right)^{\frac{1}{3}}\leqslant e_{\text{Jel}}\leqslant\zeta_{\text{BCC}}(1)\simeq-1.4442$ (17)
and conjectured that the upper bound is in fact an equality [Wigner-34, BlaLew-15, Lewin-22]. The lower bound is proved in [LieNar-75, Appendix]. The right side of (17) is the Epstein zeta function of the Body-Centered Cubic lattice $\mathcal{L}$ of unit density, containing the origin, given by the expression
@@
This is a convergent series for $\Re(s)>3$ which admits a meromorphic extension to the whole complex plane $\mathbb{C}$ with a unique pole at $s=3$. The Jellium value is then given by this extension at $s=1$ [Lewin-22, Lemma 32].
@@ 2. Proof of Theorem 3 on minimizing sequences for the liquid drop model
In this section we provide the proof of Theorem 3. The idea is to iterate several times the proof of [FraLie-15] so as to optimize the error terms.
@@
The volume $|\Omega_{n}|$ stays bounded and does not tend to 0 by assumption. Since the energy $I[\Omega_{n}]$ converges we conclude that ${\rm Per}\,(\Omega_{n})$ is bounded. By [FraLie-15, Prop. 2.1] we can find a center $a_{n}^{(1)}$ so that (after extraction of a subsequence) $\Omega_{n}-a_{n}^{(1)}\to\Omega_{*}^{(1)}\neq\emptyset$ locally with $0<m_{*}^{(1)}:=|\Omega_{*}^{(1)}|\leqslant m$ and ${\rm Per}\,(\Omega_{*}^{(1)})\leqslant\liminf_{n\to\infty}{\rm Per}\,(\Omega_{n})$. Here local convergence of sets means $L^{1}_{\rm loc}$ convergence of the corresponding characteristic functions.
@@
There is a small subtlety in the first equality in (19). From [Maggi-12, Lem. 15.12], the equality holds almost everywhere in $R_{n}^{\prime}$ and therefore we need to use that (18) in fact holds for a positive measure set of $R_{n}^{\prime}\in[R_{n},2R_{n})$.
Since the left side of (19) tends to 0 by assumption, we can use the lower semi-continuity of $I$ proved in [FraLie-15] and we arrive at
@@
We know that minimizers for $\mu_{*}$ have a finite diameter [KnuMur-13, Lemma 7.2]. We can thus find an $R>0$ such that $\Omega_{*}^{(j)}\subset B(0,R)$ for all $j=1,...,K$. From the $L^{1}$ convergence of the characteristic functions this implies
@@
$\frac{6\varepsilon_{n}}{\delta_{n}^{(j)}}\leqslant 6\varepsilon_{n}\sum_{k\neq j}\frac{1}{|a_{n}^{(k)}-a_{n}^{(j)}|}.$
@@
All the arguments above apply if the Coulomb potential $1/|x|$ is replaced by a Riesz potential $1/|x|^{s}$ with $0<s<d$ in any dimension $d\geqslant 1$, except for (14) which requires $0<s\leqslant 1$. However the finiteness of $m_{*}$ is only known for $s\leqslant 2$ [FraNam-21].

### v2

# Liquid Drop Model for Nuclear Matter in the Low Density Limit

@@ Abstract.
© 2025 by the authors. This paper may be reproduced, in its entirety, for non-commercial purposes.
@@ 1. Introduction and main result
In this paper we provide the first rigorous justification of the “gnocchi phase”, which is one of the nuclear pasta phases that are believed to form in the outer crust of neutron stars. We work with the simplest physical model used to describe this situation, called the liquid drop. It is due to Gamow [Gam29, Gam30] and was later refined by Bohr and Wheeler [BW39]. The same model can also describe microphase separation of diblock copolymers [OK86].
@@
Here ${\rm Per}\,(\Omega)$ denotes the surface area of a smooth set $\Omega$. For general measurable sets it is understood as the perimeter in the sense of geometric measure theory, see [Mag12] for a textbook introduction. The perimeter takes into account the short range interactions between the nuclei in a droplet. The second term is the electrostatic energy due to the fact that electrons and protons are, respectively, negatively and positively charged. The constant $\rho$ is the relative fraction of delocalized electrons. We have chosen units so that the local nuclear density and the constant in front of the perimeter are both equal to 1. In physical units they depend on the proton-neutron ratio. Note that protons are not able to bind without neutrons.
@@
Since $\Omega\subset\Lambda$, this implies $0\leqslant\rho\leqslant 1$ and then by symmetry we can assume $0\leqslant\rho\leqslant 1/2$.
At equilibrium the nuclear phase $\Omega\subset\Lambda$ minimizes the energy functional in (1). Working under the constraint that the system is exactly neutral, the minimum equals
@@
A natural question is to study the shape of these minimizers in the low density regime. As a first step in this direction, we will study the resulting energy asymtotics in this regime. We will first take $\Lambda\nearrow\mathbb{R}^{3}$ before we actually study the limit $\rho\to 0$. This order of the limits is important for the result to be physically meaningful.
@@ 1.1. Thermodynamic limit
Let $0\leqslant\rho\leqslant 1$. Let $\Lambda_{n}$ be a sequence of (r,L)(r,L)–Lipschitz open sets, so that
@@
For us an open set $\Lambda\subset\mathbb{R}^{3}$ is (r,L)(r,L)–Lipschitz if for any $x\in\partial\Lambda$, the set $\Lambda\cap B(x,r)$ is, in an appropriate coordinate system, equal to the epigraph of an LL–Lipschitz function. Under this assumption, one can see that the Fisher condition (5) holds for $h\leqslant Cr$. Hence (5) is only an assumption for large $h$. Note that all the conditions are satisfied if $\Lambda_{n}$ is the rescaling of a fixed (r,L)(r,L)–Lipschitz bounded open set containing the origin (for instance a cube or a ball). Our assumptions on the domain $\Lambda_{n}$ are not sharp and have been chosen because they are relatively simple.
@@
The existence of the limit (7) for the periodic model is a consequence of the seminal work [ACO09], where the optimal convergence rate was obtained as well as information about the minimizers. That this also implies that the limit in (6) exists for cubes and coincides with (7) was noted in [EFK20]. In Appendix A we provide a completely different proof of Theorem 1 and of the limit (7) using classical methods for Coulomb systems introduced by Lieb, Lebowitz and Narnhofer in the 1970s [LL72, LN75]. This proof works for any sequence $\Lambda_{n}$ satisfying the mentioned conditions.
@@
The next term in the expansion (8) describes how the gnocchis are arranged in the tomato sauce, corresponding to the low-density uniform background of electrons. The Jellium model (introduced below) tells us that the gnocchis will be placed at large distances, of order $\rho^{-1/3}$. A famous conjecture of Wigner [Wig34] states that they will in fact be regularly arranged on a Body-Centered Cubic lattice, in which case the Jellium energy is $e_{\rm Jel}\simeq-1.4442$. The interpretation of the negativity of $e_{\rm Jel}$ is that there is an effective attraction between the droplets, mediated by the background of electrons.
Figure 1. Schematic illustration of Theorem 2. For small background density $\rho$, a minimizer $\Omega$ will consist of many droplets of size one and relative distances of order $\rho^{-1/3}$.
@@
Let us mention some previous works related to Theorem 2. The first term $\mu_{*}\rho$ was derived in [KMN16] in a very dilute regime where the limits $L\to\infty$ and $\rho\to 0$ are coupled in the manner $\rho\sim 1/L^{2}$. The leading order in the regime considered here (first $L\to\infty$ and then $\rho\to 0$) was derived in [EFK20]. The ultra-dilute regime with coupled limits $\rho\sim 1/L^{3}$ was considered earlier in [CP10] for the periodic model and a two-term expansion in the same spirit as (8) was proved. This is a regime where screening effects are much less pronounced than in the situation considered in this paper.
@@
The main argument we use for proving (8) is to reduce to smaller sets of size $A\rho^{-1/3}$ with $A$ large but fixed. In any such set we expect finitely many droplets, similar to the regime studied in [CP10]. For the upper bound we involve the periodic Jellium problem defined as in (7) but for point-like particles. For the lower bound we use the Graf–Schenker inequality as in [GS95, HLS09] to remove the interactions between the different sets. We then argue by compactness to expand the energy locally in the corresponding sets of size $A\rho^{-1/3}$.
Several phase transitions are believed to occur when the density is increased from 0 to $1/2$, that is, when going deeper into the crust of the neutron star. Those are often called “pasta phases” [BW39, RPW83, HSY84, Oya93]. At small but positive $\rho$ the gnocchis are not believed to be perfect balls as they become in the limit $\rho\to 0$. They will be slightly elongated with an asymmetry increasing with $\rho$. When the density is increased beyond a certain value, instead of gnocchis one expects infinitely long cylinders called “spaghetti”, which are arranged along a two-dimensional lattice. (Just like the gnocchis may not be perfect balls, the base of the spaghetti cylinders may not be perfect disks.) When the density is further increased beyond another critical value (strictly less than $1/2$), instead of spaghettis one expects infinite slabs called “lasagna”, which are arranged along a one-dimensional lattice. In particular, at $\rho=1/2$ nucleons are expected to occupy slabs that are separated by completely unoccupied slabs of the same thickness. When $\rho$ is further increased, the story goes in the reverse way for the complement $\mathbb{R}^{d}\setminus\Omega$ until we reach an anti-gnocchi phase with $\Omega$ occupying most of $\mathbb{R}^{3}$, with holes looking likes balls at distances $(1-\rho)^{-1/3}$ and an energy behaving like for $\rho\to 0$, by symmetry. The energies in the spaghetti and lasagna regimes are given by the one- and two-dimensional analogues of the model studied here. The one-dimensional analogue of $e(\rho)$ can be computed explicitly [RW00, FL19]. The two-dimensional problem is still open and for partial results we refer to [CO07, GMS13, GMS14, Emm25]
@@ 1.3. Isolated droplets
$\boxed{\mu_{*}:=\min_{\begin{subarray}{c}\Omega\subset\mathbb{R}^{3}\\ |\Omega|>0\end{subarray}}I[\Omega],\qquad I[\Omega]:=|\Omega|^{-1}\left({\rm Per}\,(\Omega)+\frac{1}{2}\iint_{\Omega^{2}}\frac{\mathrm{d}x\,\mathrm{d}y}{|x-y|}\right).}$ (9)
@@
It is known [KM13, FL15, KMN16] that the minimum in (9) is attained for some bounded set $\Omega_{*}$, and that the problem admits no minimizer of a too large volume or too large diameter (but the infimum can be the same because we can put several copies of $\Omega_{*}$ infinitely far apart). This allows us to introduce the largest possible volume of minimizers:
@@
It is rigorously known that $m_{*}\leqslant 8$ by [FKN16] and $m_{**}\geqslant 5/2$ by [FL15, Prop. 4.1]. As an aside we mention that minimizers of $I[\Omega]$ are known to be balls if we add the constraint that $|\Omega|=m$, with given $m\leqslant 1$ [CR25].
For our proof of Theorem 2, we will need to know the rather precise behavior of minimizing sequences for the minimization problem $\mu_{*}$ in (9). The following result is an improvement of [FL15, KMN16].
@@ 1.4. Jellium model
Jellium is often associated with the name of Wigner [Wig34] who however introduced the quantum version in 1934. The classical model is also called the one-component plasma and it seems to have been proposed around 1900 by J. J. Thomson [Tho04] – based on previous ideas of W. Thomson (Lord Kelvin) – in order to describe the electrons in an atom, before the discovery of the nuclei. In this context, it is sometimes called the plum pudding model.
@@
The existence of the limit (16) was proved first by Lieb and Narnhofer in 1975 [LN75], based on a method introduced earlier by Lieb and Lebowitz in [LL72]. This model has recently reappeared in many works mentioned in the review paper [Lew22]. In particular, in a long series of works culminating in [Ser24] Serfaty and co-workers have obtained detailed results about the Jellium model with variable background density. As a byproduct, also results about the original Jellium model were obtained.
@@
$1.4508\approx-\frac{3}{5}\left(\frac{9\pi}{2}\right)^{\frac{1}{3}}\leqslant e_{\text{Jel}}\leqslant\zeta_{\text{BCC}}(1)\simeq-1.4442$ (17)
and conjectured that the upper bound is in fact an equality [Wig34, BL15, Lew22]. The lower bound is proved in [LN75, Appendix]. The right side of (17) is the Epstein zeta function of the Body-Centered Cubic lattice $\mathcal{L}$ of unit density, containing the origin, given by the expression
@@
This is a convergent series for $\Re(s)>3$ which admits a meromorphic extension to the whole complex plane $\mathbb{C}$ with a unique pole at $s=3$. The Jellium value is then given by this extension at $s=1$ [Lew22, Lemma 32].
@@ 2. Proof of Theorem 3 on minimizing sequences for the liquid drop model
In this section we provide the proof of Theorem 3. The idea is to iterate several times the proof of [FL15] so as to optimize the error terms.
@@
The volume $|\Omega_{n}|$ stays bounded and does not tend to 0 by assumption. Since the energy $I[\Omega_{n}]$ converges we conclude that ${\rm Per}\,(\Omega_{n})$ is bounded. By [FL15, Prop. 2.1] we can find a center $a_{n}^{(1)}$ so that (after extraction of a subsequence) $\Omega_{n}-a_{n}^{(1)}\to\Omega_{*}^{(1)}\neq\emptyset$ locally with $0<m_{*}^{(1)}:=|\Omega_{*}^{(1)}|\leqslant m$ and ${\rm Per}\,(\Omega_{*}^{(1)})\leqslant\liminf_{n\to\infty}{\rm Per}\,(\Omega_{n})$. Here local convergence of sets means $L^{1}_{\rm loc}$ convergence of the corresponding characteristic functions.
@@
There is a small subtlety in the first equality in (19). From [Mag12, Lem. 15.12], the equality holds almost everywhere in $R_{n}^{\prime}$ and therefore we need to use that (18) in fact holds for a positive measure set of $R_{n}^{\prime}\in[R_{n},2R_{n})$.
Since the left side of (19) tends to 0 by assumption, we can use the lower semi-continuity of $I$ proved in [FL15] and we arrive at
@@
We know that minimizers for $\mu_{*}$ have a finite diameter [KM13, Lemma 7.2]. We can thus find an $R>0$ such that $\Omega_{*}^{(j)}\subset B(0,R)$ for all j=1,,Kj=1,...,K. From the $L^{1}$ convergence of the characteristic functions this implies
@@
$\frac{6\varepsilon_{n}}{\delta_{n}^{(j)}}\leqslant 6\varepsilon_{n}\sum_{k\neq j}\frac{1}{|a_{n}^{(k)}-a_{n}^{(j)}|}.$ (22)
@@
All the arguments in the proof of Theorem 3 apply if the Coulomb potential $1/|x|$ is replaced by a Riesz potential $1/|x|^{s}$ with $0<s<d$ in any dimension $d\geqslant 1$, except for (14). The latter requires $0<s\leqslant 1$ so that the localization error of the perimeter can be controlled as in (22). Moreover, the finiteness of $m_{*}$ is only known for $s\leqslant 2$ [FN21].